please answer the following questions quickly 
 
 a) Two clindrical capacitor has radii of inner and outer plates (a,b) and (c,d) respectively. if they have same length and same capacitance then
axc=bxd how???
 
  1. Two parallel plates have charges charges Q1 and Q2 on them and the capacitance is C. The potential difference between them is   (Q1 -Q2)/2c how?
  2.  
 

Grade:11

1 Answers

om sai
8 years ago
a)capacitance per unit length is directly proportional to 1/(ln(outer radius/inner radius))    
therefore (b/a)=(d/c) 
 
b)charge on outer sides of both plates =(Q1+Q2)/2
charge on inner side of one plate is(Q1-Q2)/2  and on inner side of another plate is-(Q1-Q2)/2
 
we know that charge = (capacitance)(potential difference)
therefore potential difference =(Q1-Q2)/2C
hence proved 
these questions asked by you show that you are thorough in BASIC concepts STOP problem solving and LEARN basics FIRST